### v1

# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

Abstract
The ability of artificial intelligence agents to make optimal decisions and generalise them to different domains and tasks is compromised in complex scenarios. One way to address this issue has focused on learning efficient representations of the world and on how the actions of agents affect them, such as disentangled representations that exploit symmetries. Whereas such representations are procedurally efficient, they are based on the compression of low-level state-action transitions, which lack structural richness. To address this problem, we propose to enrich the agent’s ontology and extend the traditional conceptualisation of trajectories to provide a more nuanced view of task execution. Structurally Enriched Trajectories (SETs) extend the encoding of sequences of states and their transitions by incorporating hierarchical relations between objects, interactions and affordances. SETs are built as multi-level graphs, providing a detailed representation of the agent dynamics and a transferable functional abstraction of the task. SETs are integrated into an architecture, Structurally Enriched Trajectory Learning and Encoding (SETLE), that employs a heterogeneous graph-based memory structure of multi-level relational dependencies essential for generalisation. Using reinforcement learning as a data generation tool, we demonstrate that SETLE can support downstream tasks, enabling agents to recognise task-relevant structural patterns across diverse environments.
keywords:
Representation learning , Trajectories , Reinforcement Learning , Heterogeneous graphs , Graph Neural networks Code available here: https://github.com/CatarauCorina/setle_code
[…]
apalike
[…]
[label1] organization=Artificial Intelligence Research Centre (CitAI), Department of Computer Science, addressline=City, University of London, city=London, postcode=EC1V 0HB, country=United Kingdom
1 Introduction
Autonomous systems perceive their environment, take actions, and adapt their behaviour through experience. One key paradigm in this field is Reinforcement learning (RL) (Sutton et al.,, 1998), which enables agents to learn from interaction with their environment by using reward signals to guide decision-making. Unlike supervised learning, which relies on labelled data, or unsupervised learning, which identifies patterns in data, RL allows agents to develop policies that maximise cumulative rewards over time. This approach has led to significant advancements in Artificial Intelligence (AI), for instance in game-playing AI (Silver et al.,, 2017), robotics (Levine et al.,, 2018), and autonomous systems (Kiran et al.,, 2021). Despite these successes, RL agents face fundamental challenges when transferring knowledge from simpler environments to complex, dynamic ones requiring strong generalisation and knowledge transfer.
Representation learning plays a crucial role in addressing these challenges by developing structured, transferable encodings of an agent’s environment (Bengio et al.,, 2013; Lesort et al.,, 2018). Various approaches have been proposed, for example, hierarchical RL is particularly effective for hierarchical RL (Pateria et al.,, 2021) introduces structured abstractions for multi-step decision-making but remains constrained by predefined reward structures (Sutton et al.,, 1999; Bacon et al.,, 2017). Goal-conditioned RL enhances planning by leveraging future states as learning targets, yet still depends on explicitly defined objectives (Kaelbling,, 1993; Schaul et al.,, 2015). Disentanglement techniques refine state representations by separating controllable and uncontrollable factors, improving sample efficiency (Higgins et al.,, 2018).
Although these proposals have been instrumental in enhancing policy learning, improving sample efficiency, and facilitating task-specific generalisation, we believe that a new conceptual framework is needed for representational learning that moves beyond compression methods towards enriching structural dependencies between states and actions (Nie et al.,, 2017).
Traditional representations provide a functional foundation, however, they lack the flexibility to support generalisation and knowledge transfer effectively, which may require enriching RL’s ontology. We claim that rather than treating states as undifferentiated wholes, decomposing them into their constituent objects and integrating interaction semantics would yield a more hierarchically structured and transferable representation of task dynamics. This approach would enhance generalisation across tasks, allowing agents to recognise shared structural patterns even in novel environments.
Encoding isolated interactions between objects, however, is not enough; agents need to also learn how these interactions influence future decisions and outcomes. To model this effectively, it is essential to incorporate affordances, which describe the potential interactions available to an agent based on both its capabilities and the environment (Gibson,, 2014). For example, a chair may afford sitting to a human but not to a large animal, such as an elephant. Structured representations of affordances enable an agent to predict the effects of its interactions, providing a foundation for adaptive motor planning and effective task generalisation (Schack and Ritter,, 2013), also enable agents to anticipate how changes in context impact success or failure, reducing the risk of misgeneralisation (Khazatsky et al.,, 2021).
Despite all the advantages that representing affordances would introduce, they alone would only apply to a given state. That is, affordances do not track how interactions unfold sequentially. Tracking how interactions unfold sequentially requires maintaining a structured history of transitions, a trajectory. By retaining this structured trajectory rather than relying solely on the Markov assumption that states encapsulate all relevant history, disregarding transitions (Niv,, 2019), agents will no longer be compelled to make decisions based only on their current state and will be able to learn the dependencies that shape long-term task execution, enabling more adaptive and informed decision-making. We believe that encoding such sequences of interactions and their distant effects is essential for AI systems to adapt to novel scenarios (Eysenbach et al.,, 2022) enabling the agent to capture long-term dependencies, to avoid inefficient learning and goal misgeneralisation (Di Langosco et al.,, 2022).
We define a Structurally Enriched Trajectory (SET) as an extension of the concept of trajectory. A SET goes beyond the conventional flat state-action structure by providing a structured representation that encompasses not only sequences of state-action transitions, but also encodes critical components that influence task execution. Unlike raw trajectories, which merely record the steps taken, SETs incorporate shared low-level components (objects, interactions) and the affordances that emerge from them, encapsulating information about not just what happened but how different elements contributed to varying results.
To implement such a conceptual framework that expands the RL ontology to capture hierarchical dependencies and shared trajectory knowledge, we propose a computational model that uses hierarchical subgraphs to model interactions between states, actions, and affordances, ensuring that learnt representations retain relational dependencies crucial for task execution and generalisation, which we call Structurally Enriched Trajectory Learning and Encoding (SETLE). We construct hierarchical graphs that capture interactions between states, actions, and affordances, ensuring that learnt representations retain relational dependencies essential for task execution. These structured representations empower agents to learn compact trajectory encodings, enabling them to recognise task-relevant similarities and differences across diverse environments. Graph-based learning models provide a powerful framework for capturing hierarchical causal relationships, which we pose as essential for understanding the sequence of interactions and how actions shape long-term outcomes.
These developments present an opportunity to expand the representation of trajectories with hierarchical structures. Unlike traditional representations that treat states and actions as independent transitions, graph structures naturally encode dependencies across multiple levels of abstraction, preserving how interactions evolve over time. As the need for modelling complex dependencies increases, Graph Neural Networks (GNNs) have become a key tool for learning from structured data, enabling AI systems to capture rich relational information. Heterogeneous Graph Neural Networks (HGNNs) (Wang et al.,, 2021) extend this capability by incorporating multiple node and edge types, allowing a more detailed representation of diverse interactions in dynamic environments. Still, despite their merits, most existing graph-based approaches focus on encoding full graphs rather than leveraging subgraphs with shared components. This limits their ability to model localised structures while capturing commonalities across different subgraphs. By treating entire graphs as independent entities, these methods often fail to exploit the reusability of nodes, making generalisation more challenging. A subgraph-based approach, where fundamental components are shared, offers a more effective way to represent task dynamics and facilitate knowledge transfer.
Summarising, the contributions of this research are as follows:
  1. 1.
    A conceptual framework for RL representations that extends RL’s ontology to incorporate affordances and hierarchical trajectory structures, enabling agents to capture sequences of different states and their relational dependencies.
  2. 2.
    A formal conceptualisation of SET to expand the notion trajectories to include not only state-action sequences but also objects, interactions, affordances, and task-relevant components.
  3. 3.
    A novel architecture SETLE to extract and encode SETs
    1. (a)
      A novel Hierarchical Memory Structure: A multi-level memory system that encodes task elements at varying levels of abstraction, capturing both fine-grained interactions and high-level trajectory patterns, supporting robust generalisation across tasks.
    2. (b)
      Heterogeneous Sub-Graph Encoder: A specialised encoder designed to learn trajectory representations by leveraging subgraphs with shared components rather than treating full graphs as independent entities.
  4. 4.
    Experimental evidence suggesting that SETLE effectively captures both task-specific patterns and cross-task similarities, supporting structured trajectory learning in dynamic, episodic environments  [^1] The focus of this work is creating a representation framework. Evaluating its impact on RL’s policy learning and decision-making is beyond the scope of this work.
The remainder of this paper is structured as follows: first, we review related work on trajectory representation learning and heterogeneous graph models, contextualising the limitations of current approaches. Next, we introduce the SETLE framework, detailing its hierarchical graph construction and encoding. We then present experimental results, demonstrating SETLE’s ability to generalise across tasks. Finally, we discuss the broader implications of our findings and outline future research directions.
2 Related work
This section critically explores advances in representing the dynamic task structure in complex environments, as well as their focus, successes, and challenges. We present an analysis of outcomes and trajectories highlighting crucial shortfalls and how to address them. Finally, we examine the strengths and limitations of Heterogeneous Graph Neural Networks (HGNNs) in modelling these structured task dynamics.
Most models in representation learning have focused on compression methods which aim to reduce the dimensionality of the data while preserving relevant information from the task, for example transforming high-dimensional data into compact vector representations. In particular, representational learning in agents has been heavily aimed at encoding effective state representations, seeking to capture crucial information about the environment in a compressed and meaningful way (Wang et al.,, 2021). However, while these representations improve efficiency, they do not always generalise well across tasks, as not all aspects of a state can be directly transferred to new environments. In many cases, only certain components of the state remain relevant, while others become task-specific or redundant.
Outcomes, on the other hand, have been considered as a way to summarise the effects of an agent’s actions (Williams,, 1992). Being represented as isolated events, they fail to capture the structured dependencies and hierarchical relationships crucial for generalisation. Trajectories, in contrast, provide a richer representation by encoding the full sequence of states and actions that lead to an outcome while also preserving temporal dependencies. By maintaining the ordering and progression of interactions, trajectories enable agents to understand how decisions influence future states, supporting better generalisation across tasks.
Unlike artificial RL agents, humans do not seem to learn isolated state-action transitions, but rather develop hierarchical models of causal relationships between actions and their effects (Schack and Ritter,, 2013; Morris et al.,, 2022). This enables them to adapt their strategies dynamically, rather than relying on rigid rules or pre-defined goals. Moreover, in real-life scenarios, task dynamics are inherently context-dependent, as the same action can lead to different outcomes depending on environmental conditions (Kiesel and Hoffmann,, 2004).
A key strength of human problem-solving is the ability to use causal models to recognise structural similarities between tasks, transfer knowledge across domains, and apply past experiences to new situations with minimal relearning. For instance, riding a bike and driving a car require distinct motor skills, both involve coordinating movement, adjusting speed, and responding to environmental constraints (Wilson and Collins,, 2019). Representing outcomes plays a crucial role in shaping this learning process. The way in which a task is completed, not just whether it was successful, affects how knowledge is transferred and adapted. If an outcome aligns with prior experiences, it reinforces existing predictive models, while unexpected results prompt adjustments in strategy and refinement of learnt representations.
While outcomes intended or not provide a useful summary of performance, they do not capture the full structure of execution, including the sequential dependencies that lead to a result. Understanding a task requires not just recognising the final state but also interpreting the trajectory of actions taken to reach it, a distinction that is crucial for developing adaptive artificial systems.
A key challenge in reinforcement learning (RL) is learning structured trajectory representations rather than treating experiences as disjointed outcomes. Classically, trajectories in RL are defined as sequences of state-action pairs that evolve over time, capturing the agent’s interaction with the environment from an initial state to termination (Sutton et al.,, 1998). While this definition provides a temporal structure, it does not inherently encode hierarchical relationships beyond stepwise transitions.
One extension of this classical definition is the options framework (Precup,, 2000), which introduces temporally extended actions that allow an agent to execute multi-step behaviours before concluding. An option consists of an internal policy, an initiation set, and a termination condition, enabling an agent to select and commit to more than single-step actions. Other approaches, such as RL with trajectory feedback, shift from evaluating individual steps to assessing the entire trajectory as a whole (Efroni et al.,, 2021). This is particularly useful in scenarios where immediate feedback is sparse or impractical, such as real-world robotics or autonomous systems, where long-term dependencies must be considered. Both options and trajectory feedback offer different ways to extend decision-making beyond stepwise transitions, one by grouping actions into temporally extended choices, the other by evaluating entire trajectories, yet neither provides the level of structural detail necessary for effective generalisation and knowledge transfer. Options, as well as the trajectory feedback approach, treat elements of decision-making as indivisible units, lacking the granularity needed for structured generalisation. Options group actions into fixed temporal sequences, ignoring finer-grained execution details such as object properties, interaction dynamics, and environmental constraints. Similarly, trajectory feedback evaluates entire sequences without analysing individual contributions, making it difficult to identify structural similarities across tasks or extract reusable behavioural patterns that support generalisation and adaptation. This limitation suggests the need for more structured trajectory representations that capture not just individual actions but also high-level task structures and reusable behavioural patterns.
An approach that deals with trajectories in a more structured way is the Self-Consistent Trajectory Autoencoder (SeCTAR) (Co-Reyes et al.,, 2018). SeCTAR frames hierarchical reinforcement learning (HRL) as a trajectory-level representation learning problem. Instead of learning discrete sub-policies, SeCTAR builds continuous latent spaces of trajectories, allowing agents to represent and generate structured behaviours. While SeCTAR’s latent space encodes meaningful trajectory structures, it does not explicitly segment trajectories into transferable skills, potentially limiting its effectiveness for compositional learning and policy transfer. Methods such as the Constructing Skill Trees (CST) algorithm (Konidaris et al.,, 2011) address this limitation by automatically segmenting demonstration trajectories into reusable skills using change-point detection techniques. This approach enables agents to decompose long-horizon tasks into meaningful sub-policies, facilitating skill reuse and hierarchical planning. However, CST relies on human-provided demonstrations, which means that it does not autonomously explore or discover skills without predefined examples. A more generalisable approach would integrate structured trajectory representations with autonomous skill discovery, enabling agents to learn and refine reusable skills without requiring expert demonstrations.
Most existing trajectory representation methods in RL treat trajectories as flat sequences, focussing on linear dependencies rather than hierarchical structures. For example, the Outcome-Driven Actor-Critic (ODAC) framework (Rudner et al.,, 2021), trajectory-based prediction models (Haines et al.,, 2018), and memory-based RL frameworks such as Sequential Episodic Control (SEC) (Freire et al.,, 2024) all adopt flat representations of trajectories as a core design choice. ODAC models latent outcome representations by treating trajectories as flat sequences of state-action transitions, providing a structured yet non-hierarchical approach to inferring task success. Similarly, Haines et al., (2018) emphasises cumulative patterns in state sequences to predict outcomes, assuming linear trajectory structures to simplify trajectory modelling. Meanwhile, SEC improves adaptability and generalisation by retaining episodic state-action pairs, allowing agents to recall past interactions and improve learning efficiency in sparse-reward environments. However, SEC’s flat memory structure prevents it from capturing hierarchical trajectory dependencies, which are essential for structured learning, skill reuse, and transferability.
While these approaches enhance learning by preserving sequential dependencies, they do not inherently encode relational dependencies between entities, their interactions, and their characteristics in a task environment. Understanding how different components interact over time, how specific objects interact or how affordances shape task execution, requires a representation that can capture both temporal and relational dependencies.
Graph-based approaches provide a natural way to represent this structured knowledge, allowing the integration of sequential flow, affordance-driven interactions, and task-specific dependencies into a unified framework.
Heterogeneous Graph Neural Networks (HGNNs) offer a potential solution by using meta-paths, hops, and network schemas to capture relational dependencies between different node types (Wang et al.,, 2021). Unlike traditional graph-based methods, HGNNs can encode multi-level interactions, making them well-suited for tasks that require hierarchical reasoning. Recent work has explored how causal graph representations can further enhance trajectory-oriented learning by explicitly modelling causal dependencies between entities, ensuring that the learnt representations reflect structured transitions rather than mere correlations (Jin et al.,, 2024). Their approach constructs causal chains and applies counterfactual link prediction to refine causal reasoning in structured graphs. However, many HGNN-based causal graph learning approaches, including those in Jin et al., (2024), rely on semi-supervised learning, which assumes access to predefined causal links or labelled annotations. Their effectiveness is often limited in RL environments, where labels such as ground-truth causal effects of actions are typically unavailable or difficult to infer. This makes such methods unsuitable for agents that must learn through interaction, as they require self-supervised or adaptive learning mechanisms.
Heterogeneous Co-contrastive Learning (Wang et al.,, 2021) stands out as a self-supervised HGNN framework that overcomes these limitations by leveraging co-contrastive learning. Unlike traditional approaches, Wang et al., (2021) uses two complementary views: the network schema and the meta-path one, to train node embeddings. By contrasting these views, it captures both local and higher-order structures within the graph, providing a richer representation of static heterogeneous networks without requiring labelled data. Despite its strengths, Wang et al., (2021)’s focus on node-centric representations and static graph structures limits its applicability in dynamic, episodic environments. Such environments require representations that evolve across hierarchical and task-based episodes, reflecting the sequential and cumulative nature of actions leading to outcomes.
To enhance adaptability, Graph Transformer Co-contrastive Learning (GTC) (Sun et al.,, 2025) extends beyond node-centric HGNNs by introducing a dual-branch architecture that integrates local GNN-based aggregation with Transformer-based global modelling, allowing it to capture both fine-grained node relationships and broader contextual dependencies. Additionally, its metapath-aware Hop2Token mechanism encodes multi-hop dependencies more effectively by transforming node interactions into structured token representations, enabling the model to incorporate hierarchical information across varying graph distances. GTC outperforms traditional HGNNs in static settings by capturing high-level semantic information without succumbing to over-smoothing.
Despite recent advancements, existing methods fail to fully address the challenges of dynamic, episodic data, as they do not explicitly capture the hierarchical and cumulative nature of task dynamics. While Haines et al., (2018) and Zou and Suzuki, (2024) each contribute to different aspects of outcome learning, they lack a structured representation that models how trajectories evolve over sequences of interactions. Similarly, trajectory-based models often focus on sequence prediction, but fail to decompose trajectories into their fundamental components, such as objects and their interactions. This lack of decomposition prevents the identification of shared structural patterns across tasks.
Furthermore, graph-based approaches such as Sun et al., (2025) and Wang et al., (2021) have demonstrated the potential of structured representation learning in capturing relational dependencies between entities. However, these methods have not been designed for reinforcement learning and typically aggregate information at a global level, making it difficult to represent task-relevant interactions within individual episodes. Additionally, they do not account for shared entities, such as objects and actions, that could reoccur across different trajectories. This prevents models from recognising common patterns across tasks, limiting their ability to generalise knowledge and adapt to variations in task execution.
These limitations highlight the need to expand the ontology of reinforcement learning to include structured representations of task elements, hierarchical trajectory dependencies, and cross-episode relationships.
The next section introduces SETLE a framework that addresses these gaps by modelling trajectories as structured subgraphs, incorporating objects, actions, and affordances as key components of task execution. We first describe the model’s conceptual ontology and then its implementation.
3 Methodology
This section outlines the conceptual framework and its implementation (which can be found here https://github.com/CatarauCorina/setle_code). SETLE was designed to represent the dynamics of task execution by structuring trajectories as hierarchical subgraphs. Unlike traditional methods that aggregate information across entire graphs, we introduce a subgraph-centric approach tailored for episodic data. Each episode is treated as an independent subgraph, with its representation depending solely on its local neighbourhood.
A key distinction of our framework is the ability to handle shared nodes that may appear in multiple trajectories or episodes. These shared nodes serve as bridges between episodes, allowing the emergence of cross-episode patterns. For example:
  • 1.
    An object, such as a ball, might be used across different episodes but in varying contexts (e.g., being pushed to a goal in one task and placed in a bucket in another).
  • 2.
    Similarly, an action like push could have different effects depending on the affordances and states in multiple episodes.
By encoding these shared nodes within their respective trajectory subgraphs, the relationships learnt from one trajectory or episode can inform the understanding of similar interactions in other episodes, enhancing generalisation across tasks.
We first define the ontology of our framework then describe the process of constructing Structurally Enriched Trajectories (SETs) as heterogeneous graphs, ensuring they encode multi-level relationships between objects, interactions, states, and affordances. We then detail the methodology for learning SET embeddings, which can be used for downstream tasks, including integration into reinforcement learning pipelines.
3.1 SETLE Ontology
This section defines key components of our framework, including interactions, Structurally Enriched Trajectories (SETs), affordances, and hierarchical memory, which together provide a structured representation of task execution.
3.1.1 Interactions and Objects
Objects and interactions are structured embeddings extracted from the environment that encode task-relevant features that influence decision-making. An interaction $i_{t}$ represents the interplay between an agent-object or object-object relationship observed within the environment. However, due to the constraints of the environment used in our experiments, interactions are limited to agent-object relations, where the agent is always a ball object, interacting with various other objects in the scene.
3.1.2 Affordances
Following (Catarau-Cotutiu et al.,, 2023), an affordance is conceptualised as the learnt relationships between an (effect, reward) pair and a (concept-object, interaction) tuple. When an agent applies an action, it produces an interaction, which in turn generates an effect (a state transition) and an associated reward, both of which are stored as an affordance. This structured representation is designed to improve generalise across tasks by predicting how different interactions influence trajectories across varying contexts.
Formally, an affordance $A$ encapsulates an agent’s action, the current state, and the resulting trajectory (Catarau-Cotutiu et al.,, 2023):
$A=\{s_{t},a_{t},s_{t+1},r_{t}\}$ (1)
where $s_{t}$ is the current state, $a_{t}$ is the agent’s action, $s_{t+1}$ is the resulting state, and $r_{t}$ is the reward at time $t$.
3.1.3 Structurally Enriched Trajectories (SETs)
In reinforcement learning (RL), a trajectory is traditionally defined as a sequence of states and actions:
$\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\dots,(s_{t},a_{t}),(s_{t+1},a_{t+1}),\dots,(s_{T},a_{T})\}$ (2)
where $s_{t}\in S$ represents the agent’s state at time step $t$, $a_{t}\in A$ is the action taken, and $T$ denotes the length of the episode.
We define a Structurally Enriched Trajectory (SET) as an extension of this classical trajectory that incorporates additional structural components, capturing relational dependencies and affordance-based interactions. A SET is represented as a hierarchical graph:
$\mathcal{G}_{\tau}=(V,E,\Phi)$ (3)
Where:
$V=\bigcup_{t=0}^{T}\{O_{t},I_{t},\mathfrak{\hat{f}}_{t},s_{t}\}$ is the set of nodes, where:
$O_{t}$ represents the set of all objects present at time step $t$.
$I_{t}$ represents the set of all interactions occurring at time step $t$.
$\mathfrak{\hat{f}}_{t}$ represents the affordance at time step $t$.
$s_{t}$ is the state of the agent at time step $t$
The union across all $t$ ensures that $V$ contains the full trajectory of objects, interactions, affordances, and states throughout the episode.
$E\subseteq V\times V$
is the set of edges, encoding temporal transitions $(s_{t},\mathfrak{\hat{f}}_{t},s_{t+1})$ and structural relationships such as object interactions $(o_{i},o_{j})$.
$\Phi:V\to\mathbb{R}^{d}$
is a feature extraction function that assigns high-dimensional embeddings to objects, interactions, and states.
This enriched representation allows us to model task execution beyond stepwise transitions, capturing how task elements evolve over time and interact within different contexts.
3.1.4 Hierarchical Memory Structure
We organise multiple SETs into a hierarchical memory structure, represented as a large heterogeneous graph, that captures shared trajectory components across different episodes. The hierarchical memory is formally represented as:
$\mathcal{M}=(\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{N})$ (4)
Where:
- $\mathcal{G}_{i}$ represents an individual SET subgraph, modelling a structured trajectory of an episode.
- Each subgraph $\mathcal{G}_{i}$ can be further decomposed into a sequence of lower-level graphs $\mathcal{G}^{\prime}_{i}$, where each $\mathcal{G}^{\prime}_{i}$ corresponds to a state in the episode, decomposed into its fundamental components (objects and interactions).
- SETs share low-level components, such as objects and interactions, creating relational dependencies across multiple trajectories.
The hierarchical structure of the memory captures different levels of abstraction. At the lowest level, each state is represented as a graph $\mathcal{G}^{\prime}_{i}$, encoding the relationships between objects and their interactions. At a higher level, SETs encapsulate structured trajectories, linking sequences of state-graphs through affordances into meaningful task representations. Unlike traditional memory structures that store isolated trajectories, our hierarchical memory structure connects trajectories through shared elements, preserving both task-specific dependencies and cross-task relationships. This structured representation allows agents to retrieve relevant experiences based on shared trajectory components, improving adaptability and transfer in reinforcement learning environments.
To formally represent the hierarchical graph for a SET, we model the task environment as a Heterogeneous Information Network (HIN), defined as a graph $G=(V,E,A,R,\Phi)$.
The network schema of $G$, denoted as $\Gamma=(A,R)$, is used to describe the structure of interactions between nodes of different types, capturing the local structure. The schema is used to define direct connections among various node types (such as states, interactions, affordances and trajectories), forming the foundation for encoding relationships in heterogeneous networks.
Additionally, we define *meta-paths* to capture the composite relationships within task episodes. A meta-path $P$ is defined as a sequence $A_{1}\xrightarrow{R_{1}}A_{2}\xrightarrow{R_{2}}\dots\xrightarrow{R_{i}}A_{i+1}$, where each $A_{i}$ is a node type and each $R_{i}$ is a relation. Meta-paths describe the indirect, higher-order connections between node types, such as sequences connecting states to the enriched trajectories, which are essential for capturing complex dependencies across task hierarchies.
3.1.5 Structured Data Source: CREATE Environment
The constraints of our framework required an environment with multiple tools and manipulable elements, allowing different solution strategies within and across tasks. To ensure the capture of structured dependencies over time, tasks needed to involve multi-step execution, reinforcing hierarchical trajectory learning rather than isolated decision-making. Additionally, task goals had to vary across episodes, compelling agents to adapt their strategies dynamically instead of relying on memorized action sequences. We employ Construction, Reuse, and Extension of Actions through Transfer and Exploration (CREATE) environment (Jain et al.,, 2020), a diverse and challenging platform specifically designed to study action generalisation in reinforcement learning. It provides a range of tasks that involve dynamic interactions with various objects and tools, such as pushing a ball to a designated area, placing it inside a container, or navigating obstacles. It also offers a rich setting to demonstrate the effectiveness of our approach, with tools to explore generalisation by varying task setups across episodes, forcing agents to adapt their strategies and reuse knowledge from past experiences.
3.2 SETLE: Hierarchical Graph Construction
Hierarchical graphs were built to capture structured relationships between objects, their interactions, affordances and state transitions as SETs. We construct these representations by processing raw interactions from the environment and encoding them into heterogeneous graph nodes and edges that preserve both local and global dependencies. In this section, we cover the different operations required to obtain these graphs.
3.2.1 SET Definition Through Random Exploration
In RL, an episode consists of a complete sequence of interactions from an initial state to termination, encapsulating the trajectory an agent follows within a given task. To extract and formalise SETs (for a detailed definition, see Section 3.1.3) we employ an exploration strategy within the CREATE environment (Jain et al.,, 2020). In this setup, agents engage with tasks such as pushing a ball to a goal or placing it inside a bucket, generating diverse trajectories. To maximise variability in the trajectory space and better capture adaptive patterns and task-relevant structures, RL agents operate under a random policy rather than optimised for specific rewards.
Each episode is represented as a SET graph (see Fig.1), encapsulating the task setup, actions taken, and the result of the task, which is stored for validation and further analysis in the hierarchical memory (see 3.1.4).
An episode $E$ is defined as a sequence of state-action pairs:
$E=\{(s_{0},a_{0}),(s_{1},a_{1}),\dots,(s_{t},a_{t}),\dots(s_{T},a_{T})\}$ (5)
where $s_{t}\in S$ is the state at time $t$, and $a_{t}\in A$ is the action taken at time $t$.
A SET is labelled as success if the agent reaches the goal; otherwise, it is labelled as failure. The SET of an episode $SET(E)$ is labeled as follows:
$SET(E)=\begin{cases}\text{success},&\text{if }s_{T}\in G\\ \text{failure},&\text{otherwise}\end{cases}$ (6)
where $G$ is the set of goal states.
SETs are represented hierarchically (see Section 3.2.2) in the memory structure, including the episode’s inventory and action sequence. For each episode, the corresponding SET graph $G_{SET}$ is constructed, consisting of nodes representing states and edges representing state transitions (i.e., as seen in Fig.1).
3.2.2 Hierarchical Memory Structure
To effectively model SETs, we structured episodic data into hierarchical heterogeneous graphs, where the feature extraction functions $\Phi:V\to\mathbb{R}^{d}$ (see Section 3.1.3) assigned high-dimensional embeddings to nodes representing objects, interactions, and states. The nodes and their different relations are organised into hierarchical heterogeneous graphs with three distinct levels of abstraction: (1) a base layer for objects and interactions, where objects are extracted using the SAM model (Kirillov et al.,, 2023) and interaction dynamics are encoded via a ConvLSTM; (2) a mid-layer for states and affordances, capturing the transition dynamics and relationships between interactions and their effects; and (3) a trajectory layer for task-level abstractions, representing the cumulative impact of interactions and sequences (see Fig. 2). This hierarchical organisation is stored in a graph database and is not simply a stratification of layers but a structured representation where lower levels influence higher-level abstractions. Objects and actions influence affordances, which in turn shape state transitions and contribute to the formation of SETs. This dependency across levels enables agents to reason over task structure, capturing both local dynamics and high-level task generalisation in complex environments. Formally, the hierarchical memory graph G representing a SET consists of:
Figure 1: This illustration represents the hierarchical structure of a SET, capturing relationships between different node types in task representation. At the lowest level, interactions (light-filled circular nodes) depend on objects (darker circular nodes). Moving upward, affordances (smaller darker nodes) emerge from interactions, which then influence states (highlighted in red with distinct borders) which in turn influence the next affordance. At the highest level, the SET node connects to the structure through temporal state dependencies. The edges indicate relationships such as object dependencies, contributions, and effects. Varying line styles help differentiate these relationships. This hierarchical structure showcases how task execution is represented by linking objects, interactions, affordances, and states.
Figure 2: Hierarchical Memory building: The image shows the models used to build the hierarchical memory. The base layer extracts objects and interactions using the SAM object detection model and a pre-trained ConvLSTM, capturing fundamental task dynamics. The mid-layer represents states and affordances, modelling how interactions influence future states. At the highest level, the SET layer encodes trajectory-level abstractions, capturing long-term dependencies across sequential states. This hierarchical organisation enables efficient reasoning over task structures and adaptive decision-making.
Level 1: low-level abstractions (objects and interactions)
The lowest level of the hierarchy captures individual objects (Fig. 3) and interactions representations (Fig. 4), which form the fundamental units in the environment (Fig. 5).
Figure 3: Base-level representations: Object-level representations extracted at the base layer using the SAM object detection model. These representations capture object features in the environment and are stored in a graph database.
Figure 4: Base-level representations: Interaction-level representations extracted at the base layer using a pre-trained ConvLSTM action embedder. These embeddings capture the dynamics of object interactions and are stored in the graph database.
The relationship between objects and interactions is fundamental to model structured learning. Fig. 5 illustrates how objects interact with each other, emphasising that an interaction is defined not only by its motion dynamics but also by the entities involved in its execution.
Figure 5: Representation of two objects interacting. Interactions are characterized not only by their execution but also by the relationships they share with the objects involved. On the right, the different parameters of an Interaction are shown such as the type, and the vector representation. Representation obtained from the Neo4j (Lal,, 2015) graph database interface.
We have employed a combination of object detection and interaction representation learning techniques to construct these different representations of objects and interactions.
  1. 1.
    Object Detection for extraction of object nodes
    Objects within each episode are extracted using the SAM (Segment Anything Model) object detection model Kirillov et al., (2023), which identifies and classifies entities in the environment. A process of checking whether the object exists already is carried out using a cosine similarity function.
    $\text{Similarity}(o_{\text{new}},o_{i})=\frac{\mathbf{v}_{\text{new}}\cdot\mathbf{v}_{i}}{\|\mathbf{v}_{\text{new}}\|\|\mathbf{v}_{i}\|},\quad\forall i\in\text{Memory}$ (7)
    If $\text{If }\max_{i}\,\text{Similarity}(o_{\text{new}},o_{i})<\tau,\text{ add }o_{\text{new}}\text{ as a new node.}$. (8)
  2. 2.
    Learning Interactions Representations using ConvLSTM for the extraction of interaction nodes
    An encoder is used to model object interactions effectively, capturing their temporal and spatial dynamics within the environment. While traditional approaches like the Neural Statistician have been utilized to learn latent representations for unordered datasets by generating a posterior $q(c|D)$ over the context $c$ of each dataset $D$, their application to environments with high task similarity, such as CREATE, reveals notable limitations. Specifically, the shared context $c$ across episodes often results in poor separability between distinct tasks, hindering the generalisation of representations.
    We use a ConvLSTM-based encoder to address the challenges of differentiating between similar objects and their interactions across tasks. This encoder is designed to capture the temporal dependencies inherent in sequences of object interactions. By encoding temporal dynamics, the ConvLSTM model provides a richer and more structured representation of the task environment. Furthermore, to ensure that the embeddings effectively distinguish between similar and distinct interactions, we couple the ConvLSTM with a triplet loss function. The triplet loss enforces separation in the embedding space by bringing embeddings of similar interactions closer together while pushing embeddings of distinct ones further apart. As such we were able to produce robust and discriminative interaction representations, forming a foundational component for subsequent hierarchical encoding.
    Let $\mathcal{F}_{a}=\{f_{1},f_{2},\dots,f_{n}\mid\text{all frames associated with interaction }a\}$ represent the set of frames associated with a specific interaction $a$. For a given frame $x_{a}\in\mathcal{F}_{a}$, we denote $x_{p}\in\mathcal{F}_{a}$ as a positive sample (a frame associated with the same interaction $a$) and $x_{n}\notin\mathcal{F}_{a}$ as a negative sample (a frame associated with a different interaction).
    The triplet loss $L_{\text{triplet}}$ is defined as:
    $L_{\text{triplet}}=\max(0,d(f(x_{a}),f(x_{p}))-d(f(x_{a}),f(x_{n}))+\alpha)$ (9)
    Where $f(x)$ is the embedding function that maps a frame into the embedding space, $d(\cdot)$ is a distance function (e.g., Euclidean distance), and $\alpha$ is the margin. In our experiments, we set $\alpha=0.6$ after a hyper-parameter search of values ranging from 0.1-1.0. The final value provided optimal separation between embeddings of similar and distinct interactions.
    This margin value helps the network distinguish subtle differences between interactions while maintaining adequate separation between positive and negative pairs.
    Using this approach, we observed a clear separability in the embedding space. We applied K-means clustering to the learned embeddings and revealed the distinct clusters corresponding to different types of interactions. In the environment used (CREATE), all interactions involve a ball object interacting with various other objects. Thus, our analysis focuses on the second object, as it determines the dynamics of the interaction.
    While clustering separability is not perfect, the learned representations exhibit an emergent understanding of interaction similarity. For instance, interactions involving swings or trampolines may occasionally overlap in certain clusters, yet the model consistently groups bouncy objects together, distinguishing them from interactions involving fixed objects (Fig. 6). This behaviour highlights the network’s ability to capture underlying object properties that are essential for solving tasks in the CREATE environment.
    Figure 6: The scatter plot visualises the second object in interactions categorized as Fixed (orange, right side of the plot) or Bouncy (blue, left side of the plot), with each symbol representing a unique object (e.g. triangle, hexagon etc). The type of symbol along with their color indicates specific objects (e.g., Bouncy_Triangle, Fixed_Hexagon). The y-axis represents unique object IDs, while the x-axis categorizes objects into Bouncy or Fixed.
Level 2: Affordances and State Transitions
At the intermediate level, we represent states as heterogeneous graphs, denoted as $G_{it}^{\prime}$ (See Section 3.1.4) at time $t$. Each state comprises multiple interacting objects, connected through interaction representations that shape the decision-making process (Fig. 1). Affordances serve as critical links between states, interactions, and trajectories by capturing the transition dynamics (Fig. 7).
Affordances encode both immediate effects and long-term consequences, enabling the model to prioritise interactions that contribute to the successful completion of the task.
Figure 7: Illustration of the hierarchical structure in SETLE as extracted from the Neo4j graph database (Lal,, 2015). The second level of the hierarchy consists of states (red, from the middle node the first set of connections) connected by affordances (green, the outer-most nodes), illustrating the transitions between states via affordances.
By explicitly modelling affordance relationships, we improve the system’s ability to reason about sequential dependencies and facilitate adaptive decision-making in complex environments that move beyond traditional action-reward, capturing richer contextual information.
Level 3: High-Level Abstraction of Structurally Enriched Trajectories (SETs)
The topmost layer in the hierarchical structure corresponds to the overall SET representation. In its raw form, a SET is defined by the underlying levels, comprising states, affordances, interactions, and objects (see Section 3.1.3). However, utilising full graph structures for downstream tasks is computationally impractical. To address this, we learn a latent representations of SETs (see Section 3.3), effectively compressing the rich hierarchical information into a compact embedding. This latent representation forms the core of our proposed framework, enabling efficient and scalable utilisation in downstream tasks without losing critical task-relevant information.
The hierarchical graph structure obtained illustrated in Figure 1, captures the relationships between different levels of abstraction within an episode. The visualisation shows a diverse set of nodes, each representing a unique component of the episode, including objects, states, interactions, and affordances, linked through various relationships.
In the graph, nodes are colour-coded according to their roles:
  1. 1.
    Orange nodes represent individual objects detected in the environment, each connected by edges that denote has_object relationships to higher-level states or actions.
  2. 2.
    Pink nodes represent interaction representations
  3. 3.
    Red nodes represent states, encapsulating the conditions of the environment at specific time steps. Each state node can be further decomposed into low-level objects and interactions, creating a detailed snapshot of dynamics within each time frame.
  4. 4.
    Green nodes denote affordances, which are essential links that describe how specific actions lead to transitions from one state to another. Affordances, as connections between actions and resulting states, also encapsulate reward information, guiding the agent towards interactions that are likely to yield positive outcomes.
3.3 SETLE: SET encoder
This section details the methodology for encoding SET subgraphs, including contrastive learning techniques and triplet loss, which structure and differentiate trajectory representations to support robust learning for downstream tasks such as reinforcement learning. Building upon the structured representations introduced in Section 3.2, we now focus on encoding episodic graphs into compact latent vectors, ensuring that trajectory representations capture essential structural dependencies that facilitate effective generalisation and decision-making.
A graph encoder is trained with triplet loss to ensure that episodes with similar trajectories cluster together, distinguishing successful task executions from unsuccessful ones.
3.3.1 Collecting data for training
We collect SETs and label them to create a dataset for training, by following these steps:
  1. 1.
    Random exploration: We begin with an exploration phase where the agent performs multiple tasks within the CREATE environment. During this phase, a random policy agent operates in the environment until it reaches the goal state. For tractability, we have limited the number of steps the agent can take in an environment to 9, if it has not reached a goal state by then, the SET is labelled as unsuccessful. We repeat the process until sufficient successful trajectories have been collected for training.
  2. 2.
    SET Labelling: After each episode, we assess whether the agent successfully completed the task. If the episode ends with the agent achieving the goal state, the episode SET is labelled as: success; otherwise, it is labelled as: failure. Alongside the label, we store relevant metadata about the setup, actions taken, and objects encountered, allowing for later validation and analysis. Successful SETs are stored in the agent’s memory (see Section 3.1.4), completed with the inventory of objects, interactions, and state transitions that led to the successful completion.
Once a sufficient number of trajectories have been collected, we train the SETLE encoder using a triplet loss function.
3.3.2 Graph Encoder Architecture Based on Heterogeneous Co-contrastive Learning (HeCo)
A graph encoder is a neural architecture that transforms graph-structured data into meaningful embeddings, preserving the relationships and dependencies between entities. Unlike standard feature extractors, graph encoders operate over relational data, using the graph topology to learn structured representations.
Graph encoders rely on message passing, a core operation in Graph Neural Networks (GNNs) (Scarselli et al.,, 2008), where each node iteratively aggregates information from its neighbours to refine its representation (Gilmer et al.,, 2017). In its general form, message passing is defined as:
$H^{(l+1)}=\sigma\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}H^{(l)}W^{(l)}\right)$ (10)
Where:
  • 1.
    $H^{(l)}\in\mathbb{R}^{n\times d^{(l)}}$: Node representations at the $l$-th layer.
  • 2.
    $d^{(l)}$: Hidden dimension of the $l$-th layer.
  • 3.
    $W^{(l)}\in\mathbb{R}^{d^{(l)}\times d^{(l+1)}}$: Learnable weight matrix for the $l$-th layer.
  • 4.
    $n$: Number of nodes in the graph.
  • 5.
    $\tilde{A}=A+I$: Adjacency matrix $A$ with added self-loops (identity matrix $I$).
  • 6.
    $\tilde{D}$: Degree matrix of $\tilde{A}$.
  • 7.
    $\sigma$: Nonlinear activation function.
This process allows nodes to incorporate structural context from their neighbours, making graph encoders particularly effective for tasks that require relational reasoning.
Most existing approaches either encode entire heterogeneous graphs or focus on subgraph encoding for homogeneous data. Full-graph encoding enables global feature propagation, but it does not preserve localised dependencies. We model the hierarchical memory in SETLE as a collection of SET subgraphs, unlike standard GNNs that operate on a global graph, SETLE learns representations for independent trajectory (SET) subgraphs, where each SET node aggregates information solely from its connected states, interactions, objects, and affordances. By structuring trajectories (SETs) as subgraphs within the larger task graph (see Section 3.1.4), SETLE captures common elements, such as objects and interactions, at lower levels, allowing localised learning while maintaining global coherence across tasks.
Building on the principles of HeCo (introduced in the Related Work section), we adapt its contrastive learning approach for episodic graphs. We treat each SET as a subgraph, with the SET node acting as the central node. The central node refers to the primary node within a graph that aggregates information from surrounding nodes through both meta-path and schema-based message passing, ensuring that its embedding captures local structural dependencies as well as high-order semantic relationships within the heterogeneous graph.
The encoding operation can be formally represented as:
[…]
where $z_{SET}$ is the latent representation of a Structurally Enriched Trajectory $G_{SET}$.
To train the graph encoder, we use a triplet loss and a hybrid loss function to organise SET embeddings. Successful SETs from the same task are drawn together, while unsuccessful SETs or SETs from different tasks are pushed apart. This allows us to differentiate between successful and unsuccessful strategies, enhancing generalisation. To ensure that episodes with similar trajectories are encoded closely in the latent space, we employ a triplet loss as such:
$L_{triplet}^{SET}=\max(0,d(g(y_{a}),g(y_{p}))-d(g(y_{a}),g(y_{n}))+\alpha)$ (12)
Here:
  • 1.
    $g(y)$ is the SET embedding function.
  • 2.
    $y_{a}$, $y_{p}$, and $y_{n}$ are the anchor, positive, and negative SET samples.
  • 3.
    $d$ is a distance metric.
  • 4.
    $\alpha$ is a margin parameter.
By minimising this loss, the model learnt to align trajectory-specific embeddings, ensuring that episodes with similar trajectories (i.e., trajectories belonging to the same tasks) remained close while distinguishing those with different results.
To further improve the quality of SET representations, a modified hybrid loss function that combines the triplet loss with a cross-view contrastive loss is also used. This approach not only pushes similar SET embeddings together but also aligns two graph views: meta-path and schema, ensuring consistency across different perspectives of the episodic graph.
The two key views from HeCo are adapted to SETs as follows:
  1. 1.
    Network Schema View
    In the network schema view, we encode the structural dependencies of each episode by considering relationships between node types (e.g., state, interaction, affordance) around the central SET node. This view allows us to capture local patterns directly connected to the episode.
  2. 2.
    Metapath View
    The meta-path view captures high-order semantic relationships by encoding dependencies across sequences of connected nodes, allowing the model to learn relationships that span multiple levels of the episode’s hierarchy. The following steps outline the process of computing node embeddings under the meta-path view:
    1. (a)
      For a given node ”i” and a set of $M$ meta-paths $\{P_{1},P_{2},\dots,P_{M}\}$, which represent high-order semantic relationships (e.g., SET-St-Obj-Inter-Aff or SET-St-Aff-St):
      1. i.
        For each meta-path $P_{n}$, we apply a meta-path specific Graph Convolutional Network (GCN) to encode its characteristics:
        • A.
          Calculate the projected features $h_{i}$ and $h_{j}$ for nodes ”i” and ”j” in the meta-path.
        • B.
          Use degree information $d_{i}$ and $d_{j}$ for nodes ”i” and ”j” to compute the updated embedding $h^{P_{n}}_{i}$ based on the meta-path’s semantic similarity.
    2. (b)
      For all $M$ meta-paths, we obtain a set of embeddings $\{h^{P_{1}}_{i},\dots,h^{P_{M}}_{i}\}$ for node ”i.”
    3. (c)
      Use semantic-level attention to combine these embeddings into the final embedding $z^{mp}_{i}$ under the meta-path view:
      • i.
        Calculate the weight of each meta-path $P_{n}$ for node ”i.”
      • ii.
        Determine the importance of each meta-path in the embedding fusion.
      • iii.
        Perform a weighted sum of meta-path embeddings to produce the final node embedding.
    The selected metapaths, SET-St-Obj-Inter-Aff and SET-St-Aff-St, are designed to capture high-order semantic relationships within each episode in the RL environment.
    • (a)
      SET-St-Obj-Inter-Aff (SET - State - Object - Interaction - Affordance):
      • i.
        This meta-path reflects the hierarchical progression from the episode as a whole, through individual states, down to specific objects and actions, ultimately linking to affordances.
      • ii.
        By following this sequence, SETLE can capture the causal chain of events where each state within a SET is defined by its constituent objects and the interactions involving those objects. The inclusion of affordances in the meta-path allows the model to understand how specific object interactions influence the likelihood of transitioning to desirable or goal-oriented states.
      • iii.
        This path is crucial for learning how particular combinations of objects and their interactions contribute to successful results, enabling the model to capture the semantic importance of specific object relationships in each episode. It supports the agent in recognising patterns of interaction that are effective across similar tasks, improving object selection and result prediction in new scenarios.
    • (b)
      SET-St-Aff-St (SET - State - Affordance - State):
      • i.
        This meta-path emphasises the transitions between states within a SET, specifically focussing on how each affordance (a state-action-state link with associated rewards) facilitates state changes.
      • ii.
        By encoding this meta-path, SETLE captures the dynamics of state transitions within the trajectory. The affordance links provide a representation of the agent’s actions and the effects, highlighting which transitions are likely to lead to success or failure.
      • iii.
        This metapath is valuable for modelling temporal dependencies in SETs, as it reveals the sequences of states the agent encounters. Understanding these transitions is critical for building a representation that reflects the path to the goal, rather than just the individual states, thus helping the agent generalise across different task setups that share similar transition dynamics.
    Together, these meta-paths allow us to capture both static relationships (e.g., object interactions within a state) and dynamic relationships (e.g., state-to-state transitions driven by affordances) in the SET graph. By incorporating both aspects, we can construct a more comprehensive and semantically meaningful representation of each trajectory, allowing it to effectively generalise across tasks and improve decision-making in complex reinforcement learning environments.
Considering these two views we define the hybrid loss as:
$L_{\text{hybrid}}=\lambda L_{\text{triplet}}^{\text{SET}}+(1-\lambda)L_{\text{contrastive}}^{\text{view}}$ (13)
Here:
  • 1.
    $L_{\text{triplet}}^{\text{SET}}$ organises SET embeddings in the latent space, as previously defined:
    $L_{\text{triplet}}^{\text{SET}}=\max(0,d(g(y_{a}),g(y_{p}))-d(g(y_{a}),g(y_{n}))+\alpha)$ (14)
    where $g(y)$ is the SET embedding function, $y_{a}$, $y_{p}$, and $y_{n}$ are the anchor, positive, and negative SET samples, $d$ is a distance metric, and $\alpha$ is a margin parameter.
  • 2.
    $L_{\text{contrastive}}^{\text{view}}$ aligns the meta-path and schema view embeddings to ensure consistency across views:
    $L_{\text{contrastive}}^{\text{view}}=-\log\frac{\exp(\text{sim}(z_{\text{mp}},z_{\text{sc}})/\tau)}{\sum_{z^{\prime}\in\mathcal{Z}}\exp(\text{sim}(z_{\text{mp}},z^{\prime})/\tau)}$ (15)
    Where:
    • (a)
      $z_{\text{mp}}$ and $z_{\text{sc}}$ are the meta-path and schema view embeddings for the same SET.
    • (b)
      $\text{sim}(\cdot,\cdot)$ is a similarity function, such as cosine similarity.
    • (c)
      $\tau$ is the temperature scaling parameter.
    • (d)
      $\mathcal{Z}$ represents the set of embeddings, including positive and negative samples.
  • 3.
    $\lambda$ is a hyperparameter that controls the balance between the triplet loss and the contrastive loss, with $0\leq\lambda\leq 1$.
By combining these objectives, the hybrid loss ensures that SET embeddings are both task-specific (via the triplet loss) and consistent across graph views (via the contrastive loss).
4 Experiments
This section evaluates the effectiveness of SETLE’s hierarchical graph encoder in encoding task-specific information. The evaluation is structured into two key sets of experiments: margin sensitivity analysis and ablation studies. In the first set of experiments (S1), we assessed how different margin values in the triplet loss function affect the quality of learnt embeddings. By varying the margin parameter, we analysed its influence on the clustering of success and failure SETs in the embedding space. This experiment provided insights into the optimal margin settings that enhanced separability while maintaining robust intra-cluster cohesion.
The second set of experiments (S2) consisted of ablation studies to assess the impact of hierarchical structures and relational dependencies in SETLE’s SET design. We systematically modified or removed key components, such as hierarchical organisation, sequentiality, and affordance-based state transitions, to understand their contributions to SET representation. By comparing the clustering performance of modified versions of SET’s structure against its complete hierarchical design, we demonstrated the necessity of structured graph-based encoding for effective trajectory learning.
4.1 S1: SET Embedding Analysis: Margin and Loss Function Impact
The first set of experiments aimed to evaluate the ability of SETLE’s hierarchical graph encoder to represent enhanced trajectory information effectively. The primary focus was on assessing whether the embeddings generated by SETLE can distinguish between successful and unsuccessful episodes within and across tasks. Using episodic data collected from the CREATE environment, each episode was encoded as a heterogeneous graph (SET) that comprises objects, actions, states, and affordances.
We investigated how well SETLE’s SET embeddings captured the underlying task dynamics by analysing clustering quality using various configurations of hyper-parameters and different loss functions.
The evaluation focused on specific parameters including:
  • 1.
    Margin Values: We experimented with margin values for the triplet loss, specifically testing values of [0.1, 0.2, 0.5, 1.2, 1.5]. This parameter regulates the separability of embeddings in the latent space by controlling how far apart unsuccessful SETs are pushed from successful ones.
  • 2.
    Loss Functions: Two types of loss functions were employed:
    1. (a)
      Classical Triplet Loss: Ensures SET embeddings for similar episodes are close, while those for dissimilar SETs are further apart.
    2. (b)
      Hybrid Triplet Loss: Combines triplet loss with a cross-view contrastive loss. The contrastive component aligns embeddings from the schema and meta-path views, while the triplet component enhances task-based separability.
The following metrics were used to evaluate the separability of success and failure clusters:
  • 1.
    Silhouette Score: Measures the cohesion and separation of clusters, the values can range between -1 and 1, with higher values indicating more distinct clusters. Values closer to 1 suggest that episodes are well-clustered within their respective success or failure groups.
  • 2.
    Davies-Bouldin Index (DBI): Quantifies the average similarity ratio of each cluster with its most similar cluster, with lower values indicating better clustering performance and less overlap between clusters.
  • 3.
    Dunn Index: Assesses the ratio between the minimum inter-cluster distance and the maximum intra-cluster distance, with higher values representing greater separability and well-defined clusters.
4.1.1 Margin Impact and Results
Effect of Margin Values
We evaluated the impact of varying margin values on the clustering quality of SET embeddings. The results provided insights into the cohesion and separation of clusters for both successful and unsuccessful episodes. High cohesion within success or failure clusters indicated that the embeddings effectively captured the underlying patterns of similar SETs, while high separation between these clusters reflected the model’s ability to distinguish the dynamics of successful and unsuccessful episodes. These insights were critical for evaluating the suitability of different margin values in shaping the latent space for SET-specific encoding.
The clustering results were analysed using three standard metrics: Silhouette Score, Davies-Bouldin Index (DBI), and Dunn Index.
Table 1 summarises the clustering performance across margin values. Higher margins (1.2 and 1.5) generally yielded better separability, reflecting clearer distinctions between successful and unsuccessful episodes.
Table 1: Clustering Results for Success and Failure for SET encodings at Different Margin Values
Margin Silhouette Score (Success) Silhouette Score (Failure) DBI (Success) DBI (Failure) Dunn Index (Success) Dunn Index (Failure)
0.1 0.7632 0.8531 0.4514 0.2319 2.2497 3.4662
0.2 0.7959 0.8659 0.3205 0.2233 2.9952 3.3552
0.5 0.7943 0.8798 0.3378 0.1771 2.5417 5.3031
1.2 0.7588 0.8535 0.3603 0.2305 2.3577 3.5070
1.5 0.8137 0.8990 0.2974 0.1548 2.7778 5.5415
From the results in Table 1, it was shown that higher margin values generally improved clustering quality for both success and failure SETs. Notably:
  • 1.
    The Silhouette Score showed a consistent improvement with increasing margin, peaking at a margin of 1.5, where the success and failure clusters were well-separated.
  • 2.
    The Davies-Bouldin Index (DBI) decreased as the margin increased, reaching its lowest value at a margin of 1.5. This suggested that higher margins reduced the overlap between clusters, leading to more distinct success and failure groupings.
  • 3.
    The Dunn Index improved significantly with higher margins, especially for failure SETs, which achieved their highest separability at a margin of 1.5. This indicated that increasing the margin value results in more well-defined clusters, with greater distances between the success and failure groups.
Higher margin values proved necessary for effective clustering in this context because the CREATE environment is relatively simple, with low-resolution frames and limited visual complexity. In such environments, embeddings based on episode encodings can appear similar due to the simplicity of the scenes and fewer unique features across frames. Using higher margins (1.2 and 1.5), we ensured that embeddings for episodes with different outcomes or tasks are separated more distinctly, overcoming the limitations posed by the low frame quality and simplifying environment.
In terms of epoch loss, we did not observe significant changes with varying margin values Fig.8. The overall epoch loss progression remained relatively stable, suggesting that while higher margins did not heavily impact convergence or loss reduction during training, they did have a positive effect on the quality of clustering in the latent space. In the cluster analysis below, we examined how these higher margin values enhanced episode separability, supporting generalisation and task-based decision-making. Figure 8 illustrates the training loss for different margin values.
Figure 8: Training loss for various margin values. Higher margins improve clustering quality without significantly affecting convergence.
4.1.2 Visualization of Outcome Embeddings
To complement the quantitative analysis of clustering metrics such as Dunn Index, DBI, and Silhouette Score, we employed K-means clustering and visualisation techniques to provide a qualitative assessment of the learnt SET embeddings. By reducing the dimensionality of the embeddings using PCA, we examined the spatial distribution of clusters in two dimensions, revealing the distinctiveness of task and result representations. These visualisations offered additional insights into how well SETLE’s graph encoder captured the structural and outcome-specific features of the tasks, highlighting its ability to group similar episodes while maintaining separation between dissimilar ones. The tight grouping of points within each task cluster, along with minimal overlap between clusters of different tasks, indicated that the graph encoder has learnt meaningful representations that distinguish both the task type and the outcome.
Figure 9(a) shows the K-means clustering of SET embeddings, color-coded by task. The results highlight SETLE’s ability to capture task-level distinctions and result-specific information.
(a) Task-level clusters: K-means Clustering of SET Encodings (PCA Reduced). Each point represents a SET in the CREATE environment, color-coded by task.
(b) Outcome and task-specific clusters: Each point represents a SET in the CREATE environment, colour-coded by task, where circles mark successful SETs and crosses mark failed ones.
Figure 9: K-means clustering of SET embeddings. (a) Clustering by task only. (b) Clustering by task and outcome (success vs. failure). SETLE effectively captures both task-level distinctions and outcome-specific information.
Additionally, the separation of successful and unsuccessful SETs within the same task, indicated by distinct markers in the plot, further highlighted SETLE’s effectiveness in encoding task-specific information. For example, episodes from CreateLevelPush that successfully reached the goal state were grouped closely, while failed attempts were separated, providing evidence that SETLE captures not only the task identity but also the quality of task execution Figure 9(b).
4.1.3 Loss Functions Impact and Results
To further analyse the impact of different optimisation strategies on SET representation, we compared the performance of two loss functions: the simple triplet loss and the hybrid loss. While the triplet loss focusses on enforcing distance constraints between similar and dissimilar SET embeddings, the hybrid loss integrates contrastive learning by aligning embeddings across schema and meta-path views. This combination aims to enhance both structural consistency and task separability.
To assess the effectiveness of the hybrid loss in improving representation compared to the standard triplet loss, we visualised the SET embeddings generated by SETLE under both loss functions. Each point represents a SET, colour-coded by task, with successful episodes marked as $\square$ and unsuccessful episodes marked as $\times$. The visualisations are provided in Figures 10(a) and 10(b), where subfigure (a) shows the results with triplet loss and subfigure (b) shows the results with hybrid loss.
(a) Triplet Loss trained encoder clustering results.
(b) Hybrid Loss trained encoder clustering results.
Figure 10: Comparison of K-means clustering of SET embeddings under different loss functions. (a) Triplet loss: Provides reasonable task separation but limited inter-task alignment. (b) Hybrid loss: Enhances task separation and aligns result-specific embeddings more effectively by leveraging cross-view contrastive learning. Successful SETs ($\circ$) are closer within the same task, while failed ones ($\times$) remain distinct, reflecting improved generalization and result differentiation.
Observations on Inter-task Improvements:
The hybrid loss demonstrated inter-task improvements compared to the triplet loss:
  • 1.
    Task Separation: The hybrid loss resulted in better separation between tasks compared to the triplet loss. In Figure 10(b), tasks with shared structural dynamics (e.g., tasks using similar objects like buckets or baskets) were positioned closer in the embedding space, while maintaining distinct clusters for unrelated tasks. The hybrid loss leveraged cross-view contrastive signals (e.g., schema and meta-path views) to capture shared dynamics between tasks. This led to embeddings that better encode structural similarities and inter-task relationships, facilitating generalisation.
  • 2.
    Outcome-specific Encoding: While both loss functions distinguished successful and unsuccessful SETs within the same task, the hybrid loss brought positive and negative outcomes from the same task closer together compared to the triplet loss. This behaviour arose from the cross-view contrastive learning mechanism in the hybrid loss, which emphasised shared structural and task-specific features across both positive and negative outcomes. By leveraging complementary views (e.g., schema and meta-path), the hybrid loss encouraged the embeddings of episodes from the same task, regardless of their outcome, to cluster more closely while maintaining sufficient separation between success and failure states. This reflected a more nuanced encoding of the shared task dynamics and structural similarities inherent in both successful and unsuccessful trajectories.
The hybrid loss significantly enhanced the representation of inter-task relationships while maintaining strong intra-task and outcome-level differentiation, by using cross-view contrastive learning in addition to triplet-based alignment.
Comparison with State-of-the-Art Heterogeneous Graph Encoders
While state-of-the-art heterogeneous graph encoders such as HeCo and GTC excel in static graph settings and node-level tasks like classification and clustering, they are not directly comparable to SETLE due to differences in problem setting, data structure, and application focus. SETLE addresses the unique challenge of encoding episodic, dynamic graphs, where each episode forms a structured subgraph (SET) representing trajectories. Unlike the static, node-centric graphs typically used by HeCo and GTC, the data in SETLE comes from episodic environments like CREATE, where the graphs evolve dynamically over time and must encapsulate hierarchical abstractions across objects, interactions, states, affordances, and trajectories.
This fundamental difference in problem formulation and the nature of the data makes direct comparisons with existing heterogeneous graph encoders unsuitable and underscores the distinct contribution of SETLE to learning hierarchical SET representations.
4.1.4 Comparison on CREATE environment
In Jain et al., (2020), action representations were evaluated by encoding agent actions across episodes and analysing their separability within the embedding space. Their results highlighted effective clustering of actions, particularly for distinguishing tasks requiring distinct agent behaviours.
In contrast, SETLE shifted the focus to trajectory representations, encoding not only the actions but also the intermediate states, transitions, and affordances that cumulatively define the task’s success or failure. The comparison reveals that action representations, as emphasised by Jain et al., (2020), excel in precision for short-term predictions, such as determining the next best action in a sequence. SETLE, however, takes a broader approach by also learning interaction representations using a ConvLSTM architecture and then integrating them with state, affordance, and object information to enable comprehensive SET representation learning. By combining detailed interaction-level encoding with higher-level abstractions of task trajectories, SETLE provides a unified framework that supports both immediate decision-making and generalisation across tasks. This integration highlights the complementary nature of interaction and trajectory representations, with SETLE excelling in task-level analysis while maintaining robust interaction representation capabilities.
4.2 S2: Ablation Studies Emphasising the Model’s Design
To further investigate the importance of SETLE’s hierarchical structure, we conducted three additional experiments: random state removal, sequentiality disruption, and flattened representations. These experiments highlighted the robustness of the model and its reliance on structured representations for accurate outcome embeddings.
1. Random State Removal
Objective: Demonstrate the importance of sequential states in accurately representing SETs.
Design: Randomly removed a fixed proportion (e.g., 10%, 20%, 40%) of the states in the graph. Encode the modified SET using SETLE and compare clustering quality. Below we presented the results of removing 40% of the states, having observed that for lower values the results were insignificant.
Remarks: State removal disrupted the sequential flow and relational dependencies, leading to poorer cluster cohesion and separability (see Table 2).
Results:
Table 2: Comparison of Clustering Results for Random State Removal vs. Complete Graphs
Metric State Removal Complete Remarks
Silhouette Success 0.756 0.797 Reduced cohesion and separation of success clusters.
Silhouette Failure 0.827 0.820 Less cohesive clustering.
DBI Success 0.472 0.345 Higher overlap and less distinct success clusters.
DBI Failure 0.298 0.314 Marginally better cluster definition.
Dunn Index Success 1.838 2.561 Reduced separation and compactness of success clusters.
Dunn Index Failure 2.904 2.618 Less compact and separated failure clusters.
2. Sequentiality Disruption
Objective: Test how breaking the sequential order of states impacts SET representation.
Design: Randomly shuffled the sequence of states in each episode while keeping all nodes intact.
Remarks: Sequential disruption reduced the ability to encode cumulative effects, leading to embeddings were are less meaningful and cohesive.
Results:
  • 1.
    Silhouette Scores: Shuffled sequences exhibit lower clustering quality (success: 0.785 compared to 0.748).
  • 2.
    Dunn Index: Noticeable degradation in success cluster compactness (2.285 compared to 1.613).
  • 3.
    DBI: Increased overlap between shuffled task clusters (success: 0.378 compared to 0.494).
3. Flattened Representations
Objective: Tested the importance of hierarchical elements in trajectories by replacing the hierarchical graph with a flat graph structure.
Remarks: Flattening the graph significantly degraded the quality of the learnt task representations, highlighting the importance of relational data for meaningful task dynamics (see Table 3).
Results:
Table 3: Impact of Flattening SET Hierarchical Structure on Clustering Results
Metric Flattened Complete Impact of Flattening
Silhouette Success 0.705 0.770 Decreased cohesion and separation of success clusters.
Silhouette Failure 0.782 0.823 Reduced distinct boundaries for failure clusters.
DBI Success 0.595 0.473 Increased overlap and less distinct success clusters.
DBI Failure 0.425 0.305 Higher overlap between failure clusters.
Dunn Index Success 1.326 1.652 Reduced separation and compactness of success clusters.
Dunn Index Failure 1.864 2.888 Poorer separation and compactness of failure clusters.
4.3 Future Experiments on Complex Environments
Future work will focus on evaluating SETLE in more complex, real-world environments that involve richer dynamics, higher-dimensional state representations, and more diverse task structures. Such environments could include robotics simulations. However, accessing these environments poses significant challenges. The computational complexity of encoding high-dimensional, multi-level graphs from real-world tasks can become prohibitive, requiring both advanced hardware and efficient learning algorithms.
Despite these challenges, applying SETLE in such scenarios could provide valuable insights into its scalability and robustness, further demonstrating its ability to generalise across diverse and complex task settings. This would solidify its role as a framework capable of bridging the gap between dynamic, task-aware representations and real-world applications.
5 Discussion
In this paper, we have introduced Structurally Enriched Trajectories (SETs), a structured extension of traditional trajectory representations that captures multi-level dependencies between states, interactions, and affordances. SETs are proposed as a fundamental component of a novel computational framework for representation learning, Structurally Enriched Trajectory Learning and Encoding (SETLE), that models hierarchical task structures using a heterogeneous graph memory. By structuring trajectories as SET subgraphs, we extend HeCo’s graph embedding approach to episodic Reinforcement Learning (RL), treating each episode as a distinct subgraph and encoding entire trajectories as central nodes. Our method was able to cluster similar trajectories, constructing a task-focused latent space that can improve cross-task generalisation. Unlike traditional flat trajectory encodings, we provide a way to capture multi-level relationships using network schema and meta-path views, making it particularly effective for complex RL tasks requiring structured trajectory learning. These contributions have broader implications for representation learning, task generalisation, and lifelong AI systems, such as facilitating cross-task adaptability, refining the efficiency of learning from limited experience, and strengthening memory retention for continuous adaptation to new environments.
A key challenge in RL is to ensure that learnt representations transfer across tasks rather than remaining task-specific (D’Eramo et al.,, 2024; Chen et al.,, 2024). By incorporating hierarchical trajectory representations, we enable the recognition of shared patterns and support knowledge transfer across structurally similar settings. Hierarchical structuring is an essential mechanism for generalisation (Sancho et al.,, 2024). In human cognition, for example, learning and adaptation depend on the ability to identify and recombine meaningful substructures across experiences (Lake et al.,, 2017). Many AI models, however, struggle with compositional flexibility, limiting their adaptability to novel tasks and environments. Our framework advances an alternative approach, adopting a hierarchical graph-based representation to model task execution more effectively. This method enables agents to retain and reuse structured knowledge rather than discarding prior experiences, improving adaptation and transfer across tasks.
Although hierarchical structures have been central to reinforcement learning (RL) research for a long time (Botvinick et al.,, 2009), most current approaches depend on predefined skill segmentation. This reliance limits the ability to autonomously discover reusable structures. In contrast, we have modelled the relational dependencies between the trajectory components to ensure that abstract knowledge is connected to the execution-level details. This method aligns with broader trends in structured representation learning (Hafner et al.,, 2019; Roads and Love,, 2024), where explicitly modelling these relationships enhances transferability. By incorporating hierarchical trajectory representations into a structured memory framework, our approach offers a foundation for developing more flexible and scalable reinforcement learning models.
In addition to generalising tasks, this work also advances lifelong learning and adaptive AI. Effective open-ended learning depends on a structured memory system that enables agents to accumulate and refine knowledge over time, rather than starting over with each new task (Finn et al.,, 2017). The hierarchical memory structure in SETLE lays the groundwork for metalearning and continual learning. It allows agents to retain relevant information, adapt to new contexts, and dynamically refine their decision-making strategies. This structured approach also mitigates the issue of catastrophic forgetting, where agents can overwrite previously learnt information.
This work contributes to the broader discussion on structured intelligence in AI (Mohan et al.,, 2024; Eckstein and Collins,, 2020) by providing an effective mechanism to link low-level state-action execution with high-level conceptual reasoning. Achieving human-like adaptability requires us to move beyond mere pattern recognition; it is essential to extract, manipulate, and reuse structured representations across various domains. This exercise in human-level adaptation necessitates continuous interaction among different organisational levels of information and a well-defined hierarchical control structure (Schneider and Logan,, 2006; Hommel et al.,, 2000; Wu et al.,, 2022).
By demonstrating how graph-based trajectory modelling improves task representation, this study lays the groundwork for further exploration of structured reinforcement learning (RL) representations. Hierarchical models have played a crucial role in cognitive science and neuroscience, offering insights into how humans integrate multi-level representations to guide learning and decision-making (Eckstein and Collins,, 2020). While similar approaches in reinforcement learning have been proposed to address long-term dependencies and causal structure learning, challenges still exist in designing frameworks that can generalise beyond task-specific representations (Mohan et al.,, 2024).
While our framework offers several advantages, there are still a number of challenges that need to be addressed. For example, scalability in highly complex environments with non-linear interactions requires further research. Additionally, although hierarchical representations provide a detailed view of trajectories, their efficiency in high-dimensional tasks remains an unresolved issue. Furthermore, this work primarily focusses on representation learning, but understanding its impact on policy optimisation is a key area for future research. Moving forward, exploring how structured trajectory representations can enhance decision-making efficiency will be crucial.
Future research may extend this framework toward dynamic adaptation, enabling agents to continuously refine their representations in response to evolving task demands, pushing reinforcement learning toward lifelong, generalisable intelligence.